OpenStudy (anonymous):

A toy rocket, launched from the ground, rises vertically with an acceleration of 18 m/s^2 for 16 s until its motor stops. Disregarding any air resistance, what maximum height above the ground will the rocket achieve? The acceleration of gravity is 9.8 m/s^2. Answer in km.

OpenStudy (anonymous):

I did y = vi +1/2at^2 where y is the height, and plugged it in like this y = 0 + 1/2(18)(6)^2 y = 324 meters, then converted it to km, .324... but it's wrong, do I need to incorporate gravity somewhere in there?

OpenStudy (anonymous):

Love, your taking your values wrong S= ut+1/2gt*2 (s=Height) 1/2x18x16*2=2304m 2304m=2.3km There you go (:

OpenStudy (anonymous):

Jfc, I feel so dumb. Thank you, although it still says 2.304 is incorrect although I don't know how/why.

OpenStudy (anonymous):

Could u send me the link?

OpenStudy (anonymous):

It probably won't let you open it: https://quest.cns.utexas.edu/student/images/problem_pdf?courseuserassignment=20660548 but it's the last problem (16)

OpenStudy (anonymous):

What is the answer? If i know, then I would be able to figure out the method

OpenStudy (anonymous):

It won't tell me until the problem closes at midnight Monday, so unfortunately I cannot tell you now. But you've been so much help already, so thank you anyway:)

OpenStudy (anonymous):

You're welcome (: